# The Central Role of the Loss Function in Reinforcement Learning

### v2

@@ 4.5 Second-Order Bounds for Online RL with Optimistic MLE: Benefits of Distributional RL
$\textstyle\sigma^{2}(p_{h}(x_{h},a_{h}))\leq 2e\sigma^{2}(Z^{\pi}_{h}(x_{h},a_{h}))+H\delta_{\operatorname{dis}}^{\operatorname{RL}}(p,\pi).$
This implies the corollary:
$\mathbb{E}_{\pi}[\sigma^{2}(p_{h}(x_{h},a_{h}))]\lesssim\sigma^{2}(Z^{\pi})+H^{2}\delta_{\operatorname{dis}}(p,\pi).$
$\mathbb{E}_{\pi}[\sigma^{2}(p_{h}(x_{h},a_{h}))]\lesssim\mathbb{E}_{\pi}[\sigma^{2}(c_{h}+V^{\pi}_{h+1}(x_{h+1}))]+H\delta_{\operatorname{dis}}(p,\pi).$
Proof of Lem. 16.
Fix any $p,\pi$. We use the shorthand $\delta_{t}(x,a)=\delta_{t}^{\operatorname{dis}}(p,\pi,x,a)$ to simplify notation. First, note that the corollary follows from the main claim since the law of total variance (LTV) implies $\mathbb{E}[\sigma^{2}(Z^{\pi}_{h}(x_{h},a_{h}))]\leq\sigma^{2}(Z^{\pi})$.
Lemma 17 (LTV).
For any random variables $X,Y$:
$\textstyle\sigma^{2}(Y)=\mathbb{E}[\sigma^{2}(Y\mid X)]+\sigma^{2}(\mathbb{E}[Y\mid X]).$
We now establish the main claim.
Step 1. We first show the following claim by induction: for all $h$,
$\textstyle\sigma^{2}(p_{h}(x_{h},a_{h}))\leq$ $\textstyle\sum_{t=h}^{H}(1+\frac{1}{H})^{t-h}\mathbb{E}_{\pi}[\,8H\delta_{t}(x_{t},a_{t})$
$\textstyle 2\sigma^{2}(c_{t}+\bar{p}_{t+1}(x_{t+1},\pi(x_{t+1})))\mid x_{h},a_{h}]$ (24)
The base case $h=H+1$ is true since $\sigma^{2}(p_{H+1})=0$. For the induction step, fix any $h\in[H]$ and suppose that the induction hypothesis (IH; Eq. 24) is true for $h+1$.
By our second-order lemma for variance (Eq. 10),
$\textstyle\sigma^{2}(p_{h}(x_{h},a_{h}))\leq$ $\textstyle(1+\frac{1}{H})\sigma^{2}(\mathcal{T}^{\operatorname{D},\pi}_{h}p_{h+1}(x_{h},a_{h}))$
$+8H\delta_{h}(x_{h},a_{h}).$
Then, we use LTV to condition on $c_{h},x_{h+1}$ (i.e., the outer mean/variance are w.r.t. $c_{h},x_{h+1}$, the inner mean/variance are w.r.t. $p_{h+1}$): $\sigma^{2}(\mathcal{T}^{\operatorname{D},\pi}_{h}p_{h+1}(x_{h},a_{h}))$ is equal to
$\textstyle\mathbb{E}[\sigma^{2}(p_{h+1}(x_{h+1},\pi(x_{h+1}))\mid c_{h},x_{h+1})]$
$\qquad\textstyle+\sigma^{2}(c_{h}+\bar{p}_{h+1}(x_{h+1},\pi(x_{h+1}))).$
We bound the first term by the IH, which completes the proof for Eq. 24.
Step 2. By the above claim and $(1+\frac{1}{H})^{H}\leq e$, we have
$\textstyle\sigma^{2}$ $\textstyle(p_{h}(x_{h},a_{h}))\leq 8H\delta_{\operatorname{dis}}(p,\pi)+$
$\textstyle 2e\sum_{t=h}^{H}\mathbb{E}_{\pi}[\sigma^{2}(c_{t}+\bar{p}_{t+1}(x_{t+1},\pi(x_{t+1})))\mid x_{h},a_{h}].$
Step 3. Lastly, it suffices to convert the above variance term to $\sigma^{2}(c_{t}+V^{\pi}_{t+1}(x_{t+1}))$, since $\sigma^{2}(Z^{\pi}_{h}(x_{h},a_{h}))=\sum_{t=h}^{H}\mathbb{E}_{\pi}[\sigma^{2}(c_{t}+V^{\pi}_{t+1}(x_{t+1}))\mid x_{h},a_{h}]$ by LTV. To perform this switch in variance, observe that:
$\left\lvert\bar{p}_{h}(x_{h},\pi(x_{h}))-V^{\pi}(x_{h})\right\rvert\lesssim\sum_{t=h}^{H}\mathbb{E}_{\pi}[\sqrt{\delta_{t}(x_{t},a_{t})}],$ (25)
by the PDL followed by Lem. 1. Thus, since $\sigma^{2}(X)\leq 2\sigma^{2}(Y)+\sigma^{2}(X-Y)$ (i.e., $\sigma(\cdot)$ satisfies triangle inequality), we have
$\textstyle\sigma^{2}(c_{t}+\bar{p}_{t+1}(x_{t+1},\pi(x_{t+1})))$
$\textstyle\leq 2\sigma^{2}(c_{t}+V^{\pi}_{t+1}(x_{t+1}))$
$\textstyle+2\sigma^{2}(\bar{p}_{t+1}(x_{t+1},\pi(x_{t+1}))-V^{\pi}_{t+1}(x_{t+1}))$
$\textstyle\leq 2\sigma^{2}(c_{t}+V^{\pi}_{t+1}(x_{t+1}))+H\sum_{t=h}^{H}\mathbb{E}_{\pi}[\delta_{t}(x_{t},a_{t})],$
where the last inequality used Eq. 25 and Cauchy-Schwarz. This we have shown that
$\textstyle\sigma^{2}$ $\textstyle(p_{h}(x_{h},a_{h}))\lesssim H^{2}\delta_{\operatorname{dis}}(p,\pi)+$
$\textstyle 4e\sum_{t=h}^{H}\mathbb{E}_{\pi}[\sigma^{2}(c_{t}+\bar{p}_{t+1}(x_{t+1},\pi(x_{t+1})))\mid x_{h},a_{h}].$
By Lem. 16, we can bound Eq. 23 by
@@
By Lem. 15 and the pigeonhole principle, the error terms $\mathcal{E}_{\operatorname{mle}}^{\operatorname{RL}}$ can be bounded similarly as before: we can bound $\sum_{k=1}^{K}\delta_{\operatorname{dis}}^{\operatorname{RL}}(p^{k},\pi^{k})$ by $\widetilde{\mathcal{O}}(d_{\operatorname{mle}}H\beta)$ if UA is false, and by $\widetilde{\mathcal{O}}(Ad_{\operatorname{mle}}^{\operatorname{V}}H\beta)$ if UA is true. This finishes the proof of Thm. 9. ∎
4.6 Solving Offline RL with Pessimism
@@
Since we may take $\pi=\pi_{f}$, this is technically stronger than Assump. 3. Nevertheless, Assump. 6 is also satisfied in low-rank MDPs by the linear function class $\mathcal{F}^{\operatorname{lin}}$ and so changing from Assump. 3 to Assump. 6 does not change any conclusions we make.
As a historical remark, Alg. 4 was first proposed with the squared loss $\ell_{\operatorname{sq}}$ under the name BCP by [69] and then extended with the mle loss $\ell_{\operatorname{mle}}$ under the name P-DISCO by [61].
@@
$C^{\widetilde{\pi}}:=\max_{h\in[H]}\max_{x,a}\frac{d^{\widetilde{\pi}}_{h}(x,a)}{\nu_{h}(x,a)}.$
@@
Under Assump. 6, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 4 with $\beta=2\ln(H|\mathcal{F}|/\delta)$ has the following guarantees each loss function:
@@
We only prove the bce case as the squared loss case follows essentially the same structure. The key difference in the proof, compared to online RL, is that we use pessimism instead of optimism.
Lemma 18 (Pessimism).
Let $\ell=\ell_{\operatorname{bce}}$. Under Assump. 3, for any $\delta\in(0,1)$, setting $\beta=\Theta(\ln(H|\mathcal{F}|/\delta))$. Then, w.p.a.l. $1-\delta$, for all $\pi\in\Pi$, (a) $\mathcal{E}^{\operatorname{RL}}_{\operatorname{bce}}(\hat{f}^{\pi},\nu)\leq\frac{2H\beta}{n}$, and (b) $\min_{a}\hat{f}^{\pi}_{1}(x_{1},a)\geq V^{\pi}$.
Proof of Lem. 18.
The proof is nearly identical to that of Lem. 12 where we show that w.p.a.l. $1-\delta$, (1) all elements of the version space have low excess risk and (2) $Q^{\pi}$ lies in the version space. The only difference is that $\hat{f}^{\pi}$ is defined as the argmax rather than argmin, so that we have pessimism (greater than $V^{\pi}$) instead of optimism. ∎
By Lem. 18, we have $V^{\widehat{\pi}}-V^{\widetilde{\pi}}\leq\min_{a}f_{1}^{\widehat{\pi}}(x_{1},a)-V^{\widetilde{\pi}}$. Then, by definition of $\widehat{\pi}$, we further bound this by $\min_{a}f_{1}^{\widetilde{\pi}}(x_{1},a)-V^{\widetilde{\pi}}$. Now, we decompose with PDL:
@@
By importance sampling and Lem. 18, the error terms can be bounded by $\widetilde{\mathcal{O}}(C^{\widetilde{\pi}}\cdot\frac{H\beta}{n})$. This completes the proof of Thm. 10. ∎
Second-Order Guarantees for Offline DistRL.
We now show that DistRL with the mle loss can yield second-order guarantees, recovering the main results of [62]. We make a few minor changes to the pessimistic offline RL algorithm. First, the policy class is the set of greedy policies w.r.t. the means of the conditional distribution $\Pi_{\mathcal{P}}=\{\pi_{\bar{p}}:p\in\mathcal{P}\}$. Second, we use the
For offline DistRL, we use the policy-wise DistBC.
@@
Under Assump. 7, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 5 with $\beta=2\ln(H|\mathcal{P}|/\delta)$ has the following guarantee: for any comparator policy $\widetilde{\pi}\in\Pi_{\mathcal{P}}$,
@@
Since $V^{\widetilde{\pi}}\leq\sigma^{2}(\widetilde{\pi})$, this implies a first-order bound as well. This variance bound can be much tighter in near-deterministic settings where the comparator’s variance is near zero, but its cost is far from zero. However, as was the case in online RL, DistRL still has the drawbacks of requiring a distributional class and DistBC. While these are more stringent conditions in theory, DistRL has achieved state-of-the-art in many offline RL tasks as well [45], suggesting that the benefits of DistRL can outweight the stronger modeling assumptions in practice. The proof of Thm. 11, which we omit due to space, follows from the same argument as the proof of Thm. 10, coupled with the variance arguments from Thm. 9. The interested reader may find the full proof in [62].
4.7 Fitted Q-Iteration Algorithms for hybrid RL
@@
In this section, we discuss a potential solution that pivots to the hybrid RL setting, where the learner is both given an offline dataset $\mathcal{D}^{\operatorname{off}}$ and can interact with the environment [55]. We show that Fitted-Q Iteration (FQI) [50], a computationally efficient algorithm, can also enjoy first- and second-order guarantees by simply regressing with the bce and mle losses. Intuitively, the offline dataset mitigates the need for optimism, while the online interactions mitigate the need for pessimism. Thus, by assuming access to both offline and online RL data, we can bypass the need for optimizing over version spaces.
As a historical note, the FQI algorithm in the hybrid setting was first proposed with the squared loss $\ell_{\operatorname{sq}}$ under the name Hy-Q by [55]. Our extensions to the bce and mle losses are novel. As in offline RL, we use $C^{\widetilde{\pi}}$ to denote the coverage coefficient of the comparator policy $\widetilde{\pi}$ under the data generating distribution of $\mathcal{D}^{\operatorname{off}}$. We also assume the offline dataset to be as large as the number of interactions, i.e., $|\mathcal{D}^{\operatorname{off}}|\geq\Omega(K)$.
@@
Under Assump. 3 and $|\mathcal{D}^{\operatorname{off}}|\geq\Omega(K)$, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 6 has the following guarantees for each loss function:
@@
    $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\widetilde{\pi}}\leq\widetilde{\mathcal{O}}(H\sqrt{K\cdot(d+C^{\widetilde{\pi}})\beta}),$
@@
    $\textstyle\sum_{k=1}^{K}V^{\pi^{k}}-V^{\widetilde{\pi}}\leq\widetilde{\mathcal{O}}\big{(}H\sqrt{{\color[rgb]{1,0,0}V^{\widetilde{\pi}}K}\cdot(d+C^{\widetilde{\pi}})\beta}$
@@
Indeed, the hybrid RL bounds contain both the structural condition such as eluder dimension and the coverage coefficient $V^{\widetilde{\pi}}$. This form will be made clear in the proof, which simply combines the prior online and offline RL results. We also highlight that [7] analyzed FQI with $\ell_{\operatorname{bce}}$ in the pure offline setting and proved a first-order bound that depends on a stringent global coverage coefficient $C^{\Pi}=\max_{\widetilde{\pi}\in\Pi}C^{\widetilde{\pi}}$, which is needed to analyze FQI in the pure offline setting [14].
@@
$\textstyle\sum_{k=1}^{K}V^{\pi^{k}}-V^{\widetilde{\pi}}=$ $\textstyle\sum_{k=1}^{K}\mathbb{E}[V^{\pi^{k}}-\min_{a}f^{k}_{1}(x_{1},a)]$
@@
Under Assump. 5 and $|\mathcal{D}^{\operatorname{off}}|\geq\Omega(K)$, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 7 has the following guarantee: for any comparator policy $\widetilde{\pi}\in\Pi_{\mathcal{P}}$,
$\textstyle\sum_{k=1}^{K}V^{\pi^{k}}-V^{\widetilde{\pi}}\leq\widetilde{\mathcal{O}}\big{(}H^{2.5}(d+C^{\widetilde{\pi}})\beta$
@@ 5 Discussion and Conclusion
From the one-step CSC to online, offline and hybrid RL, we see time and time again that the loss function plays a central role in the adaptivity and efficiency of decision making algorithms. The classical squared loss always converges at a slow $\widetilde{\mathcal{O}}(1/\sqrt{n})$ rate and cannot adapt to easier problem instances with heteroskedasticity. The bce loss can serve as a drop-in improvement that yields first-order bounds with a much faster $\widetilde{\mathcal{O}}(1/n)$ rate when the optimal cost is small. Switching from conditional-mean learning to conditional-distribution learning, the mle loss can tighten the bounds further with a second-order guarantee, that is bounds that converge at a $\widetilde{\mathcal{O}}(1/n)$ rate in near-deterministic settings even if the optimal cost is large. Crucially, these gaps in performance are not merely theoretical as they have been observed many times by the deep RL community [19, 10, 29, 7, 45]. The theory outlined herein is very general and can be applied to a wide range of settings including imitation learning [21], model-based [24, 63], risk-sensitive RL [60, 64], and robust bandits [36], and RL [12]. Moreover, the principles herein can lead to better algorithms for real-world applications such as pre-training and post-training large language models [27, 1, 65], learning query optimizers [40, 66] and many more. We hope to have not only clearly demonstrated that the loss function choice is important in RL, but also to inspire the reader to seek out opportunities for better loss functions to improve their decision-making algorithms.
6 Appendix
In the appendix, we prove the pigeonhole lemma for eluder dimension (Lem. 9).
Lemma 19.
Let $E:=\sup_{p\in\mathcal{P},\psi\in\Psi}\left\lvert\mathbb{E}_{p}\psi\right\rvert$. Fix any $N\in\mathbb{N}$, $\psi^{(1)},\dots,\psi^{(N)}\in\Psi$, and $p^{(1)},\dots,p^{(N)}\in\mathcal{P}$. Let $\beta$ be a constant s.t. $\sum_{i<j}|\mathbb{E}_{p^{(i)}}\psi^{(j)}|^{q}\leq\beta^{q}$ for all $j\in[N]$. Then, $\sum_{j=1}^{N}\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert\leq\inf_{\varepsilon_{0}\in(0,1)}\{N\varepsilon_{0}+\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})\cdot(2E+\beta^{q}\ln(E\varepsilon_{0}^{-1}))\}$.
Proof of Lem. 19.
Fix any $q\in\mathbb{N}$; the proof will be for the $\ell_{q}$ eluder dimension. We say a distribution $\nu\in\mathcal{P}$ is $\varepsilon$-independent of a subset $\Gamma\subset\mathcal{P}$ if there exists $\psi\in\Psi$ s.t. $\left\lvert\mathbb{E}_{\nu}\psi\right\rvert>\varepsilon$ but also $\sum_{p\in\Gamma}^{n}(\mathbb{E}_{p}\psi)^{q}\leq\varepsilon^{q}$. Conversely, we say $\nu$ is $\varepsilon$-dependent on $\Gamma$ if for all $\psi\in\Psi$, we have $\left\lvert\mathbb{E}_{\nu}\psi\right\rvert\leq\varepsilon$ or $\sum_{p\in\Gamma}^{n}(\mathbb{E}_{p}\psi)^{q}>\varepsilon^{q}$. For any $\Gamma\subset\mathcal{P}$ and $\nu\in\mathcal{P}$, we let $N(\nu,\Gamma,\varepsilon_{0})$ denote the largest number of disjoint subsets of $\Gamma$ that $\nu$ is $\varepsilon_{0}$-dependent on. We also use the shorthand $p^{(<j)}=\{p^{(1)},p^{(2)},\dots,p^{(j-1)}\}$.
Claim 1: If $\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert>\varepsilon$, then $N(p^{(j)},p^{(<j)},\varepsilon)\leq\beta^{q}\varepsilon^{-1}$. By definition of $N:=N(p^{(j)},p^{(<j)},\varepsilon)$, there are disjoint subsets $S^{(1)},\dots,S^{(N)}\subset p^{(<j)}$ s.t. each $S^{(i)}$ satisfies $\sum_{p\in S^{(i)}}\lvert\mathbb{E}_{p}\psi^{(j)}\rvert>\varepsilon$ since $\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert>\varepsilon$ by premise. Thus, summing over all such subsets gives $N\varepsilon<\sum_{i<j}\lvert\mathbb{E}_{p^{(i)}}\psi^{(j)}\rvert^{q}\leq\beta^{q}$, proving Claim 1.
Claim 2 (Pigeonhole): For any $\varepsilon_{0}$ and any sequence $p^{(1)},\dots,p^{(N)}\in\mathcal{P}$, there exists $j\leq N$ such that $N(p^{(j)},p^{(<j)},\varepsilon_{0})\geq\lfloor\frac{(N-1)}{\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})}\rfloor$. Recall that if $p^{(1)},\dots,p^{(L)}\subset\mathcal{P}$ satisfies for all $j\in[L]$, $p^{(j)}$ is $\varepsilon_{0}$-independent of $p^{(<j)}$, then $L\leq\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})$ by definition. To prove the claim, we maintain $J:=\lfloor\frac{(k-1)}{\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})}\rfloor$ disjoint sequences $S^{(1)},\dots,S^{(J)}\subset p^{(<k)}$ s.t. each $S^{(i)}$ has the property that each element is $\varepsilon$-independent of its precedessors. We initialize $S^{(1)}=\cdots=S^{(J)}=\emptyset$ and iteratively add elements $p^{(1)},\dots,p^{(N)}$ until $p^{(j)}$ is $\varepsilon_{0}$-dependent on all these disjoint subsequences, at which point the claim is proven. If there exists a subsequence which $p^{(j)}$ is $\varepsilon_{0}$-independent of, we add $p^{(j)}$ to that subsequence, which preserves the invariant condition. This process indeed terminates since otherwise one subsequence would have more elements than $\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})$, a contradiction.
Claim 3: For any $\varepsilon$, $\sum_{j=1}^{N}\mathbb{I}[\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert>\varepsilon]\leq(\beta^{q}\varepsilon^{-1}+1)\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon)+1$. Let $\kappa$ denote the left hand sum and so let $i_{1},\dots,i_{\kappa}$ be all the indices $j$ s.t. $\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert>\varepsilon$. By Claim 2, there exists $j\leq\kappa$ s.t. $\lfloor\frac{(\kappa-1)}{\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon)}\rfloor\leq L(p^{(i_{j})},p^{(<i_{j})},\varepsilon)$. Then by Claim 1, this is further upper bounded by $\beta^{q}\varepsilon^{-1}$. Rearranging proves the claim.
Concluding the proof. For any $\varepsilon_{0}$, we have
$\textstyle\sum_{j=1}^{N}\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert=\textstyle\sum_{j=1}^{N}\int_{0}^{E}\mathbb{I}[\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert>y]\mathrm{d}y$
$\leq\textstyle N\varepsilon_{0}+\sum_{j=1}^{N}\int_{\varepsilon_{0}}^{E}\mathbb{I}[\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert>y]\mathrm{d}y$
$\overset{(i)}{\leq}\textstyle N\varepsilon_{0}+\int_{\varepsilon_{0}}^{E}\{(\beta^{q}y^{-1}+1)\operatorname{EluDim}_{q}(\Psi,\mathcal{P},y)+1\}\mathrm{d}y$
$\overset{(ii)}{\leq}\textstyle N\varepsilon_{0}+\int_{\varepsilon_{0}}^{E}\{(\beta^{q}y^{-1}+1)\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})+1\}\mathrm{d}y$
$\overset{(iii)}{\leq}\textstyle N\varepsilon_{0}+\operatorname{EluDim}(\Psi,\mathcal{P},\varepsilon_{0})(2E+\beta^{q}\ln(C\varepsilon_{0}^{-1})),$
where (i) is by Claim 3, (ii) is by monotonicity of the eluder dimension, and (iii) is by $\int_{\varepsilon_{0}}^{E}y^{-1}=\ln(E\varepsilon_{0}^{-1})$. ∎

### v3

@@ 6.2 Second-Order Bounds with MLE
$\textstyle\sigma^{2}(p_{h}(x_{h},a_{h}))\leq 2e\sigma^{2}(Z^{\pi}_{h}(x_{h},a_{h}))+H\delta_{\operatorname{dis}}^{\operatorname{RL}}(p,\pi),$
where $\sigma^{2}(Z^{\pi}_{h}(x_{h},a_{h}))$ denotes the variance of the random variable $Z^{\pi}_{h}(x_{h},a_{h})$. This also implies the corollary:
$\mathbb{E}_{\pi}[\sigma^{2}(p_{h}(x_{h},a_{h}))]\lesssim\sigma^{2}(Z^{\pi})+H^{2}\delta_{\operatorname{dis}}^{\operatorname{RL}}(p,\pi).$
By Sec. 6.2, we can bound Eq. 19 by
@@
By Sec. 6.2 and the pigeonhole principle, the error terms $\mathcal{E}_{\operatorname{mle}}^{\operatorname{RL}}$ can be bounded similarly as before: we can bound $\sum_{k=1}^{K}\delta_{\operatorname{dis}}^{\operatorname{RL}}(p^{k},\pi^{k})$ by $\widetilde{\mathcal{O}}(d_{\operatorname{mle}}H\beta)$ if UA is false, and by $\widetilde{\mathcal{O}}(Ad_{\operatorname{mle}}^{\operatorname{V}}H\beta)$ if UA is true. This finishes the proof of Thm. 9. ∎
7 Offline RL via Pessimistic Regression
@@
Since we may take $\pi=\pi_{f}$, this is technically stronger than Asm. 3. Nevertheless, Asm. 6 is also satisfied in low-rank MDPs by the linear function class $\mathcal{F}^{\operatorname{lin}}$ and so changing from Asm. 3 to Asm. 6 does not change any conclusions we make. As a historical remark, Alg. 4 was first proposed with the squared loss $\ell_{\operatorname{sq}}$ under the name BCP by [73] and then extended with the mle loss $\ell_{\operatorname{mle}}$ under the name P-DISCO by [64].
@@
$C^{\widetilde{\pi}}:=\max_{h\in[H]}\max_{x,a}\frac{d^{\widetilde{\pi}}_{h}(x,a)}{\nu_{h}(x,a)}.$ (21)
@@
Under Asm. 6, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 4 with $\beta=2\ln(H|\mathcal{F}|/\delta)$ has the following guarantees each loss function:
@@
We only prove the bce case as the squared loss case follows essentially the same structure. The key difference compared to online RL is that we establish pessimism instead of optimism. {restatable}[Pessimism]lemmaBCERLPessimism Let $\ell=\ell_{\operatorname{bce}}$. Under Asm. 3, for any $\delta\in(0,1)$, setting $\beta=\Theta(\ln(H|\mathcal{F}|/\delta))$. Then, w.p.a.l. $1-\delta$, for all $\pi\in\Pi$, (a) $\mathcal{E}^{\operatorname{RL}}_{\operatorname{bce}}(\hat{f}^{\pi},\nu)\leq\frac{2H\beta}{n}$, and (b) $\min_{a}\hat{f}^{\pi}_{1}(x_{1},a)\geq V^{\pi}$.
Proof of Sec. 7.
The proof is essentially identical to that of Sec. 6.1 where we show that w.p.a.l. $1-\delta$, (1) all elements of the version space have low excess risk and (2) $Q^{\pi}$ lies in the version space. The only difference is that $\hat{f}^{\pi}$ is defined as the argmax rather than argmin, so that we have pessimism (greater than $V^{\pi}$) instead of optimism. ∎
By Sec. 7, we have $V^{\widehat{\pi}}-V^{\widetilde{\pi}}\leq\min_{a}f_{1}^{\widehat{\pi}}(x_{1},a)-V^{\widetilde{\pi}}$. Then, by definition of $\widehat{\pi}$, we further bound this by $\min_{a}f_{1}^{\widetilde{\pi}}(x_{1},a)-V^{\widetilde{\pi}}$. Now, we decompose with PDL:
@@
By importance sampling and Sec. 7, the error terms can be bounded by $\widetilde{\mathcal{O}}(C^{\widetilde{\pi}}\cdot\frac{H\beta}{n})$. This completes the proof of Thm. 10. ∎
7.1 Second-Order Bounds via Distributional RL
We now show that DistRL with the mle loss can yield second-order guarantees, recovering the results of [65]. We make a few minor changes to the pessimistic offline RL algorithm. We consider the set of greedy policies w.r.t. the means of the conditional distribution as our policy class $\Pi_{\mathcal{P}}=\{\pi_{\bar{p}}:p\in\mathcal{P}\}$. Following Xie et al. [73], Wang et al. [65], in offline RL, we posit the policy-wise distributional BC condition.
@@ 7.1 Second-Order Bounds via Distributional RL
Under Asm. 7, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 5 with $\beta=2\ln(H|\mathcal{P}|/\delta)$ has the following guarantee: for any comparator policy $\widetilde{\pi}\in\Pi_{\mathcal{P}}$,
@@
Since $\sigma^{2}(\widetilde{\pi})\leq V^{\widetilde{\pi}}$, this implies a first-order bound as well. This variance bound can be much tighter in near-deterministic settings where the comparator’s variance is near zero, but its cost is far from zero. However, as was the case in online RL, DistRL still has the drawbacks of requiring a distributional class and DistBC. While these are more stringent conditions in theory, DistRL has achieved state-of-the-art in many offline RL tasks as well [47], suggesting that the benefits of DistRL can outweight the stronger modeling assumptions in practice. The proof of Thm. 11, which we omit due to space, follows from the same argument as the proof of Thm. 10, coupled with the variance arguments from Thm. 9. The interested reader may find the full proof in [65].
8 Computational Efficiency via Hybrid RL
@@
In this section, we discuss a solution via the hybrid RL setting where the learner can access an offline dataset $\mathcal{D}^{\operatorname{off}}$ with good coverage and also interact with the environment. We show that Fitted-Q Iteration (FQI) [53], a computationally efficient algorithm, can also enjoy first- and second-order guarantees by simply regressing with the bce and mle losses. The FQI algorithm in the hybrid setting was first proposed with the squared loss $\ell_{\operatorname{sq}}$ by [58] and our extensions to the bce and mle losses are novel.
Intuitively, the offline dataset mitigates the need for optimism, while the online interactions mitigate the need for pessimism – together, they obviate the need for maintaining a version space. For the following guarantees, we use $C^{\widetilde{\pi}}$ to denote the coverage coefficient of the comparator policy $\widetilde{\pi}$ under the data generating distribution of $\mathcal{D}^{\operatorname{off}}$. We also assume the offline dataset to be as large as the number of interactions, i.e., $|\mathcal{D}^{\operatorname{off}}|\geq\Omega(K)$ [58].
@@
Under Asm. 3 and $|\mathcal{D}^{\operatorname{off}}|\geq\Omega(K)$, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 6 has the following guarantees for each loss function:
@@
    $\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\widetilde{\pi}})\leq\widetilde{\mathcal{O}}(H\sqrt{K\cdot(d+C^{\widetilde{\pi}})\beta}),$
@@
    $\textstyle\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\widetilde{\pi}})\leq\widetilde{\mathcal{O}}\big{(}H\sqrt{{\color[rgb]{1,0,0}V^{\widetilde{\pi}}K}\cdot(d+C^{\widetilde{\pi}})\beta}$
@@
Indeed, the hybrid RL bounds contain both the structural condition such as eluder dimension and the coverage coefficient $V^{\widetilde{\pi}}$. This form will be made clear in the proof, which simply combines the prior online and offline RL results. We finally discuss some related works. [7] analyzed FQI with $\ell_{\operatorname{bce}}$ in the pure offline setting and proved a first-order bound that depends on the much larger global coverage coefficient $C^{\Pi}=\max_{\widetilde{\pi}\in\Pi}C^{\widetilde{\pi}}$, which is needed to analyze FQI in the pure offline setting [15]. Also, [49] is able to achieve computationally efficient learning low-rank MDPs without requiring offline data with good partial coverage; however, their bounds are neither first nor second-order. It would be interesting future work to adapt the techniques in this paper to derive variance dependent bounds for computationally efficient algorithms without requiring good offline data.
@@
$\textstyle\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\widetilde{\pi}})=$ $\textstyle\sum_{k=1}^{K}\mathbb{E}[V^{\pi^{k}}-\min_{a}f^{k}_{1}(x_{1},a)]$
@@
Under Asm. 5 and $|\mathcal{D}^{\operatorname{off}}|\geq\Omega(K)$, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 7 has the following guarantee: for any comparator policy $\widetilde{\pi}\in\Pi_{\mathcal{P}}$,
$\textstyle\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\widetilde{\pi}})\leq\widetilde{\mathcal{O}}\big{(}H^{2.5}(d+C^{\widetilde{\pi}})\beta$
@@ 9 Discussion and Conclusion
From the one-step CSC to online, offline and hybrid RL, we see time and time again that the loss function plays a central role in the adaptivity and efficiency of decision making algorithms. The classical squared loss always converges at a slow $\widetilde{\mathcal{O}}(1/\sqrt{n})$ rate and cannot adapt to easier problem instances with heteroskedasticity. The bce loss can serve as a drop-in improvement that yields first-order bounds with a much faster $\widetilde{\mathcal{O}}(1/n)$ rate when the optimal cost is small. Switching from conditional-mean learning to conditional-distribution learning, the mle loss can tighten the bounds further with a second-order guarantee, that is bounds that converge at a $\widetilde{\mathcal{O}}(1/n)$ rate in near-deterministic settings even if the optimal cost is large. Crucially, these gaps in performance are not merely theoretical as they have been observed many times by the deep RL community [21, 11, 31, 7, 47]. The theory outlined herein is very general and can be applied to a wide range of settings including imitation learning [23], model-based [26, 66], risk-sensitive RL [63, 67], and robust bandits [38], and RL [13]. Moreover, the principles herein can improve algorithms for post-training large language models [29, 1, 68, 77], learning query optimizers [42, 69] and many more real-world applications. We hope to have not only clearly demonstrated that the loss function choice is important in RL, but also to inspire the reader to seek out opportunities for better loss functions to improve their decision-making algorithms.
10 Acknowledgments
This material is based upon work supported by a Google PhD Fellowship and grants NSF IIS-1846210, NSF IIS-2154711, NSF CAREER 2339395 and DARPA Learning Network Cyberagents (LANCER). We also thank the editor and anonymous reviewers for useful discussions and feedback.